OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 4 to 9. If the perimeter of the larger rectangle is 54 inches, what is the perimeter of the smaller rectangle?

OpenStudy (wasiqss):

24 , medals requied :P

OpenStudy (wasiqss):

54*4/9

jhonyy9 (jhonyy9):

4/9 =x/54 54*4/9 =x 6*4=x 24=x